OpenStudy (anonymous):

Convert the polar coordinates below to Cartesian coordinates. Give exact answers. (1, 7pi/6)

OpenStudy (anonymous):

I'm not sure why, but I cannot solve this one. the answers I originally got were -.5 and -.87, but both were incorrect.

OpenStudy (whpalmer4):

\[\theta =\frac{7\pi}{6}\] is the diagonally opposite point on the unit circle from \[\theta = \frac{\pi}{6}\]so we can just flip the signs on the \((x,y)\) after finding it in more familiar territory: 

OpenStudy (ikram002p):

cool :) r=1=x^2+y^2 tan theta=y/x

OpenStudy (whpalmer4):

\[x = -r \sin \frac{\pi}{6} = -(1)(\frac{1}{2}) = -\frac{1}{2}\]\[y = -r\cos \frac{\pi}{6} = -1\frac{\sqrt{3}}{2} = -\frac{\sqrt{3}}{2}\] I do note that the problem says exact answers, which your decimal isn't for the second part.

OpenStudy (whpalmer4):

but it appears to me that you understand the concept correctly. hard to get too worked up about an error that doesn't reflect a lack of understanding of the underlying problem.

OpenStudy (anonymous):

Thank you. Turns out my homework didn't like the decimal form!